OpenStudy (anonymous):

How to prove that at x=1 |x-1| is not differentiable?

OpenStudy (anonymous):

is that y=

OpenStudy (anonymous):

Yes! y=|x-1|

OpenStudy (anonymous):

y=x-1 and y=-(x-1) from the difintion of the absolute value

OpenStudy (anonymous):

Is this the derivative? f'(x)= { x-1, x>1 1-x, x<1

OpenStudy (anonymous):

If so, how do I use the right and left hand derivatives to show that the derivative DNE?

OpenStudy (anonymous):

the right derivative =1 the left derivative =-1

hartnn (hartnn):

do you need to use limits here ?

OpenStudy (anonymous):

Yes!

hartnn (hartnn):

differentiable if \(\large \lim \limits_{h \to 0} \dfrac{f(c+h)-f(c)}{h}\) exists

hartnn (hartnn):

so, take the left hand limit for this our 'c' = 1 for left hand limit, you will take the function of x>1 or x< 1 ?

OpenStudy (anonymous):

Is it x<1?

hartnn (hartnn):

yes, so our f(x) = 1-x f(c+h) =... ?

hartnn (hartnn):

f(c) = f(1) = 0 in both cases

OpenStudy (anonymous):

Oh ok I see! Can I write Lim x--> 0 f'(x)?

hartnn (hartnn):

Lim x--> 0 f'(x) for what ?

hartnn (hartnn):

f'(x) itself is not defined...

OpenStudy (anonymous):

I see..how would you do this without newton's quotient and using chain rule instead?

hartnn (hartnn):

i am guessing, f(x) = |x-1| f'(x) = (|x-1|)' d/dx (x-1) but since d/dx |x| does not exist , (|x-1|)' also DNE so, f'(x) DNE

OpenStudy (anonymous):

Isn't the derivative of |x-1| x-1, x>1 1-x, x<1

OpenStudy (anonymous):

I think?

hartnn (hartnn):

the "definition " of |x| is x when x>0 -x when x<0 so what u said is the definition, not the derivative

OpenStudy (anonymous):

Oops! I meant 1, x>1 -1, x<1?

hartnn (hartnn):

thats correct

hartnn (hartnn):

since they are not = at x=1 derivative DNE

hartnn (hartnn):

but we didn't use chain rule here

OpenStudy (anonymous):

Is chain rule Lim x--> c+ f'(x)?

hartnn (hartnn):

what i know of chain rule is \([f(g)]' = f'(g)\times g'\)

OpenStudy (anonymous):

Sorry, bad terminology, I meant using derivatives, not newton's!

hartnn (hartnn):

yeah, we used derivatie and proved that it didn't exist, as left hand derivative (-1) not = to right hand derivative (+1), derivative at x=1 DNE

OpenStudy (anonymous):

Ok I get it! Thank you!

hartnn (hartnn):

welcome ^_^